value: 1.00 points Problem 5-4 Calculating Annuity Present Values [LO An investment offers $7,600 per year for 16 years, with the first payment occurring one year from now. Assume the required return is 9 percent. What is the value of the investment today? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Present value What would the value be if the payments occurred for 41 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Present value What would the value be if the payments occurred for 76 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Present value What would the value be if the payments occurred forever? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Present value Hints eBook & Resources Hint#1 Check my work
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

1.Present value=7600[1-(1.09)^-16]/0.09

=7600*8.312558193

=$63175.44(Approx).

2.Present value=7600[1-(1.09)^-41]/0.09

=7600*10.78656899

=$81977.92(Approx).

3.Present value=7600[1-(1.09)^-76]/0.09

=7600*11.09521307

=$84323.62(Approx).

4.Present value=Annual cash flows/required rate

=7600/0.09

=$84,444.44(Approx).